# Consumption Partial Insurance

### v4

@@ 5.3.2 Results by age and position in the income distribution
Table 4 reports results from the quadratic specification (see table C.10 for the linear specification). $\phi^{(1)}$ decreases with age, so the average permanent shock impacts consumption less as households age. In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle. $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age. It drops from $0.006$ to $-0.008$ as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket. As a result, large negative permanent shocks impact consumption more severely at older ages. Consider, again, $\zeta_{it}=-0.5$. At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.$=0.080$), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.$=0.077$).
@@
Guvenen et al. (2021) document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the young. In our data, skewness of permanent shocks becomes more negative as households age; it more than triples in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9). The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption technology itself changes ($\phi^{(2)}$ more negative). The type of events that underlie large income downturns at older ages – such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption technology that differs between the old and the young. This does not mean that moderate bad shocks are worse insured among the old; for small shocks, the decreasing pattern in $\phi^{(1)}$ more than counteracts the increasing magnitude of $|\phi^{(2)}|$.
@@
We next consider how partial insurance changes with the household’s position in the income distribution before the shocks hit at $t$, i.e., in the distribution of $y_{it-1}$. Our linear income process and first-difference approach does not directly link the transmission parameters to the levels of income, unlike ABB whose nonlinear income process naturally allows this. We thus take an empirical approach and estimate the model over different cuts of the distribution of $y_{it-1}$. Sample size considerations prevent us from being overly granular; we tried estimation over deciles, quartiles, and tertiles, and the latter was the most robust. For brevity we report results in the bottom (‘poor’) and top (‘rich’) tertiles. [^41] We target the same moments as earlier, broken down by tertiles of the distribution of past income $y_{it-1}$. The moments of shocks are also income level-specific (appendix table C.9).
@@
At face value, $\phi^{(2)}$ is more negative among the poor, suggesting that a bad shock of a given size transmits more strongly among them. While factually correct given the estimates, this provides an incomplete view of the underlying dynamics because the distribution of income risk is vastly different between poor and rich. Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but very negative among the rich, with a coefficient of $-4.86$ (multiple times larger than baseline skewness, and highly significant). One may thus study how large bad/good shocks from a given position in the underlying distribution, for example shocks in the 3rd and 97th percentiles, affect consumption in each case. Simulations are needed to infer the specific shocks. [^42] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9. We provide details on similar simulations in section 6. Adopting this approach, the bad shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) but it shifts $c_{it}$ by $-0.035$, twice as much, among the rich (pass-through $0.084$). By contrast, the good shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but shifts $c_{it}$ by only $0.006$ among the rich (pass-through $0.057$). These findings agree with ABB, who simulate the transmission of 10- and 90-percentile shocks and also find a stronger response to bad news among households in the top end of the income distribution and a stronger response to good news among households in the bottom end. The reported $\Delta c_{it}$ are also numerically in the same ballpark (impulse responses with heterogeneity in their appendix). Our results thus suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
@@
Our approach highlights two distinct forces at play in the response of consumption to income shocks: the underlying nature of income risk (i.e., the income parameters) and the consumption technology (i.e., the consumption parameters). The larger response of the rich to shocks in the bottom of the distribution is not because their consumption technology differs vastly from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates the response to said shocks), but mostly due to stronger downwards risk that the rich face. To infer that, one needs to evaluate the marginal distribution of income risk. ABB’s quantile method can flexibly estimate the consumption response over income quantiles and compare responses to, say, 10- and 90-percentile shocks, but it does not measure the higher-order features of the distribution these shocks come from. This may lead inadvertently to comparisons between consumption responses to seemingly different shocks, when these shocks may in fact be very similar to one another since the underlying distribution is skewed and leptokurtic. For instance, our estimates for skewness and kurtosis reveal that the 10- and 90-percentile shocks are both negative and have almost the same size among the poor; this is why we looked further out at the 3rd and 97th percentiles. Technically, in the context of a linear income process, and assuming transitory shocks away and no interaction between the income level and its shock, ABB estimate
@@
where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ is a uniform shock denoting the quantile of the shocks’ distribution $F_{t}$. They then simulate the consumption pass-through over different values for $y_{it-1}$ and $u_{it}$, without measuring $F_{t}$. Our model allows us to evaluate the second and higher moments of $F_{t}$ and estimate pass-through rates that explicitly shift with the sign and size of the economically relevant shock, i.e., with $\zeta_{it}=F_{t}^{-1}(u_{it})$. It can thus describe more transparently how income shocks of different signs/sizes shift consumption.
How does a less negative $\phi^{(2)}$ among the rich reconcile with a more negative $\phi^{(2)}$ among the old, who are presumably also rich? The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between the nature of income risk and the consumption technology. Skewness of permanent shocks is twice as large among the rich ($-4.86$) as it is among the old ($-2.39$). The consumption distribution is not too different between the two, however, so the rich must have more insurance to a bad shock of given size. This explains why $\phi^{(2)}$ is four times smaller among them. In both cases (old, rich), bad permanent shocks from a given position in the distribution decrease consumption more strongly than among their counterparts (young, poor).
@@ 5.3.3 Results by wealth and education
We next consider four subsamples formed on the basis of household wealth and education. We use conventional definitions for ‘low’ versus ‘high’ wealth (the cutoff is median wealth in the sample) and ‘no’ versus ‘some’ college (the cutoff is 12 years of education). [^43] Wealth is defined as the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt. Appendix table C.1 provides summary statistics for these subsamples.
@@
Table 6 reports results from the quadratic model (see table C.12 for the linear model). Three points stand out. First, $\phi^{(1)}$ decreases monotonically as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college). As $\phi^{(1)}$ measures the transmission of the average permanent shock, these groups possess gradually more insurance to average shocks; this reflects the insurance role of wealth, which increases monotonically across the groups. Second, $\phi^{(2)}$ is negative but increases (i.e., less negative) as we move across groups, with the exception of group 2 (low wealth and some college) where it equals zero. Negative shocks thus mostly transmit more strongly than positive shocks, but wealthier and more educated people have more insurance to bad shocks of given size. These people face, however, much stronger downwards income risk. This pattern is consistent with the gradient between poor and rich in table 5. Third, the transmission parameters of transitory shocks are mostly statistically insignificant, with two exceptions. In group 1 (low wealth, no college), $\psi^{(1)}$ is significant but negative, which is counterintuitive at first, yet $\psi^{(2)}$ is also negative, large, and significant. Large negative transitory shocks thus reduce consumption substantially among low wealth, low education households. In group 3 (high wealth, no college), $\psi^{(2)}$ is positive and significant, suggesting that large good shocks have a larger pass-through than smaller ones. We argued earlier that this helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in studies that employ covariance restrictions. Note that the size of group 3 is by far the smallest in our sample.
@@
5.3.4 Link with structural model
@@
Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, and income distributions. Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
@@
Madera (2019) finds that earnings and nondurable consumption are correlated in the middle of the earnings distribution, but less so in the tails compared to earnings and durable consumption. Negative shocks are associated with larger pass-through rates than equal-sized positive shocks, albeit at twice the magnitude that we estimate here.
@@
De Nardi et al. (2019) find that tail risk has large welfare costs because, among other things, it increases precautionary savings (thus also enables more insurance to a given shock). They show that the BPP method that relies on second-order moments implies a substantially higher pass-through rate than the true rate in the model.
@@
Busch and Ludwig (2024) simulate a model with borrowing constraints. We do not model such constraints, yet a strength of our approach is that it enables us to measure partial insurance without taking a stance on the precise economic mechanisms at play. They find that the marginal propensity to consume increases with the severity of a negative permanent shock as households move towards the constraint, while it decreases with the magnitude of a positive shock. Their pass-through rates are slightly higher than ours, though their model does not feature other insurance over self-insurance.
@@
Guvenen et al. (2024) simulate a model with retirement, stochastic lifetimes, and altruism. The consumption response varies with the size and sign of shocks; the pass-through of the average negative permanent shock is between 0.29 and 0.46, depending on age, higher than ours but lower than BPP. They show that measuring pass-through as in BPP overestimates the true rate by 1.5 times. Without non-employment shocks, agents are willing to give up almost about 22.3% of consumption to live in a world without risk.
@@
We read these results as not too different from ours. Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow, so those shocks end up affecting consumption by relatively more than other shocks. The magnitude of our pass-through rates seems uniformly lower than theirs (to some extent this is because these models focus on self-insurance), though other quasi reduced-form studies using the recent PSID, most notably ABB, find similar rates to ours.
6 Welfare costs of tail income risk
In this section, we calculate the welfare costs of income risk, which, as we found previously, is negatively skewed, highly leptokurtic, and features a nonlinear pass-through to consumption. We are particularly interested in contrasting the implications of linear and nonlinear transmission. To do so, we draw a sequence of income shocks over the lifecycle, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare. Our interest lies in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2). We assume $\mu_{1}<0$ for identification. Appendix table D.1 reports the parameter estimates and the (excellent) model fit. In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it},$
These reflect the model-derived linear and quadratic consumption functions (4) and (5), with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data. We simulate consumption over a working life of 35 years, starting at age 30. In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households. [^44] We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample.
In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively. The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of linear transmission. Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is much higher relative to a Gaussian density.
Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shocks they experience, in the case of $\gamma=2$. The first row presents the average cost across all households, unconditional on a given shock. This is close to zero across all three scenarios, or even negative in some cases, suggesting that households may in fact prefer tail over Gaussian risk. Inspecting the parameter estimates in table D.1, tail risk is associated with a very small probability (about 5%) that shocks come from a distribution with negative mean and large variance, and a very high probability that they come from a distribution with small positive mean but tiny variance. As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the variance of Gaussian risk. For many, Gaussian risk in fact implies larger variability, so households are on average willing to pay to move to the non-Gaussian environment.
Table 7: Welfare costs of non-Gaussian relative to Gaussian income risk
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
All households 100% -0.81 -0.88 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
The next rows zoom in on households who experience large negative permanent shocks at least once in their life. These are households who happen to draw from the first distribution in the mixture and the consequences of tail risk are clearly felt among them. Those who face a permanent shock at least $3\sigma$ below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.8% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly. This increases to 6.36% and 8.54% for shocks respectively $4\sigma$ and $5\sigma$ below the mean – all shown in column 1. These costs are already sizable, and they increase further by about 60% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.23%) for those who experience the worst shocks. Those costs would be even larger if we considered a counterfactual zero risk scenario as in Guvenen et al. (2024).
These welfare calculations are ‘back-of-envelope’, as they neglect heterogeneity in income risk and consumption transmission, they treat wealth as exogenous, and they do not account for borrowing constraints. Appendix table D.2 repeats the calculations allowing for age-specific risk and consumption transmission; the costs are in the same ballpark as the baseline. Appendix table D.3 calculates the welfare costs by wealth; the least wealthy have costs almost twice as large as the baseline, while the wealthier have smaller costs, as expected. As the likelihood that a borrowing constraint binds depends on wealth or age, the age- and wealth-specific results are indirectly reflective also of such constraints. Without a structural model, we cannot allow households to switch wealth groups endogenously (e.g., in response to an income shock), yet the baseline costs lie in the middle of the wealth-specific costs, so exogenous wealth is unlikely to have caused a serious bias in our calculations. The welfare estimates are also robust to small variations in the choice of $\gamma$ (appendix table D.4).

### v5

@@ 5.3.2 Results by age and position in the income distribution
Table 4 reports results from the quadratic specification (see table C.10 for the linear specification). $\phi^{(1)}$ decreases with age, so the average permanent shock affects consumption less as households age. In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle. $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age. It drops from $0.006$ to $-0.008$ as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket. As a result, large negative permanent shocks impact consumption more severely at older ages. Consider, again, $\zeta_{it}=-0.5$. At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.$=0.080$), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.$=0.077$).
@@
In our data, skewness of permanent shocks becomes more severe as households age, increasing in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9). This is consistent with Guvenen et al. (2021), who document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the young. The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption response itself changes ($\phi^{(2)}$ becomes more negative). The type of events that underlie large income downturns at older ages, such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption response that differs between the old and the young. This does not mean that moderate bad shocks are worse insured among the old; for small or moderate shocks, the decreasing pattern in $\phi^{(1)}$ offsets the increasing magnitude of $|\phi^{(2)}|$.
@@
Next, we consider how partial insurance varies over the household’s position in the income distribution prior to the realization of shocks, i.e., the distribution of $y_{it-1}$. Our linear income process and first-difference approach does not directly link the transmission parameters to the levels of income, therefore we take an empirical approach and estimate the model over different quantiles of $y_{it-1}$. Sample size considerations prevent us from being overly granular; we tried estimation over deciles, quartiles, and tertiles, and the latter was the most robust. For brevity we report results in the bottom (‘poor’) and top (‘rich’) tertiles. [^43] We target the same moments as earlier, broken down by tertiles of the distribution of past income $y_{it-1}$. The moments of shocks are also income level-specific; we report them in appendix table C.9.
@@
At face value, $\phi^{(2)}$ is more negative among the poor, and highly significant, suggesting that a given bad shock transmits more strongly among them. However, this provides only a partial picture of the underlying dynamics because the nature of income risk is vastly different between poor and rich. Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but negative among the rich, with a coefficient of $-4.86$ (multiple times larger than the baseline, and highly significant). One may thus study how large bad/good shocks from a given position in the underlying distributions, e.g., shocks in the 3rd or 97th percentiles, affect consumption in each case. Simulations are needed to infer the value of the specific shocks. [^44] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9. We provide details on similar simulations in appendix D.
Adopting this approach, the lower-tail shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) and by $-0.035$, twice as much, among the rich (pass-through $0.084$). By contrast, the upper-tail shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but by only $0.006$ among the rich (pass-through $0.057$). In other words, households in the top of the income distribution (the rich) exhibit a stronger consumption change in response to bad news, while households in the bottom (the poor) exhibit a stronger consumption change in response to good news.
These findings agree with ABB, who simulate the transmission of shocks in the 10th and 90th percentiles. The reported $\Delta c_{it}$ – impulse responses with heterogeneity in their appendix figure S30 – are in the same ballpark, especially for bad shocks. While this is reassuring, our findings further suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels. Our approach also highlights two distinct forces at play: the underlying nature of income risk (i.e., income parameters) and the nature of the consumption response (i.e., consumption parameters). The larger consumption change among the rich after a lower-tail shock is not because their consumption response differs materially from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates said shocks), but mostly due to stronger downside risk that they face (the magnitude of the bad shock is simply larger in this case).
@@
To distinguish between these two forces, one must evaluate the marginal distribution of income risk. While ABB’s quantile method can flexibly estimate consumption changes over different income quantiles – and compare changes after, say, 10- and 90-percentile shocks – it does so without explicitly accounting for the higher-order features of the distribution these shocks come from. Technically, in the context of a linear income process, and assuming transitory shocks away and no interaction between the income level and its shock, ABB estimate
@@
where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ denotes the quantile of the shocks’ distribution $F_{t}$. They then simulate the consumption pass-through over different values of $y_{it-1}$ and $u_{it}$, without accounting for $F_{t}$. This may lead inadvertently to comparisons between consumption responses to very different income shocks, even if those shocks come from the same position in their respective distributions (the nature of income risk varies over the distribution of income). [^45] Apparently, this may also lead inadvertently to comparisons between consumption responses to seemingly different shocks, when these shocks may in fact be similar to one another. For instance, our estimates for skewness and kurtosis reveal that the 10- and 90-percentile shocks are both negative and have almost the same size among the poor; this is why we looked farther out at the 3rd and 97th percentiles. By contrast, our approach allows us to evaluate the second and higher moments of $F_{t}$ and estimate pass-through rates that explicitly shift with the economically relevant shock, i.e., with $\zeta_{it}=F_{t}^{-1}(u_{it})$. It can thus describe more transparently how income shocks of different signs/sizes shift consumption.
How does a less negative $\phi^{(2)}$ among the rich (smaller response to larger bad shocks) reconcile with a more negative $\phi^{(2)}$ among the old (stronger response), who are presumably also rich? The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between income risk and consumption response. Skewness of permanent shocks is twice as large among the rich ($-4.86$) as among the old ($-2.39$). The consumption distribution is similar for both, so the rich must have overall more insurance to a given bad shock. This explains why $\phi^{(2)}$’s magnitude is four times smaller among them. In both cases (old, rich), bad permanent shocks from a given position in the distribution reduce consumption more than among their counterparts (young, poor).
@@ 5.3.3 Results by wealth and education
We next consider four subsamples formed on the basis of household wealth and education. We use conventional definitions for ‘low’ versus ‘high’ wealth (the cutoff is median wealth in the sample) and ‘no’ versus ‘some’ college (the cutoff is 12 years of education). Appendix table C.1 provides summary statistics for these subsamples. [^46] Wealth is defined as the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt.
@@
Table 6 reports results from the quadratic model (see table C.12 for the linear model). Three points stand out. First, $\phi^{(1)}$ decreases monotonically from group 1 (low wealth, no college) to group 4 (high wealth, some college). $\phi^{(1)}$ measures the transmission of the average permanent shock, so these groups possess gradually more insurance to average shocks. This reflects the insurance role of wealth, which increases monotonically across groups.
Second, $\phi^{(2)}$ is negative but increases (i.e., less negative) as we move rightwards across groups, with the exception of group 2 (low wealth, some college) where it equals zero. Negative shocks thus transmit weakly more strongly than positive shocks among everyone, but wealthier and more educated agents have more insurance to bad shocks of a given size. They also face, however, much stronger downside income risk, a pattern consistent with the gradient in income risk between poor and rich in the earlier discussion.
Third, the transmission parameters of transitory shocks are mostly insignificant, with two exceptions. In group 1 (low wealth, no college), $\psi^{(1)}$ and $\psi^{(2)}$ are both significant and negative, leading to the odd prediction that larger bad shocks may increase consumption slightly while smaller good shocks reduce it. In group 3 (high wealth, no college; the smallest group in our sample), $\psi^{(2)}$ is positive and significant, suggesting that large good shocks have a larger pass-through than smaller ones (thus opposite from group 1). Both cases are at odds with the bulk of estimates for the transmission of transitory shocks throughout the paper that suggest that such shocks are fully insured, so these results remain a puzzle.
@@
5.3.4 Link with structural model and some estimates of welfare costs
@@
Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, or income distribution. Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
@@
Madera (2019) finds that earnings and nondurables are correlated mainly in the middle of the earnings distribution, with weaker correlations in the tails – especially when compared to earnings and durables. This suggests that tail shocks pass through to durables (which we do not model) more than to nondurables. Negative shocks exhibit larger pass-through rates than equal-sized positive shocks, albeit at twice the magnitude that we estimate here.
@@
De Nardi et al. (2019) find that tail risk has large welfare costs because, among other things, it increases precautionary savings enabling more insurance to a given shock. They show that the BPP method relying on the workhorse consumption-income covariance implies a substantially higher pass-through rate than the true rate in the model.
@@
Busch and Ludwig (2024) simulate a model with borrowing constraints. While we do not explicitly model such constraints, a strength of our approach is that it measures insurance without relying on specific economic mechanisms. They find that the marginal propensity to consume rises with the size of a negative permanent shock as agents approach the constraint, but falls with the size of a positive shock. Their pass-through rates are slightly higher than ours, though their model excludes other form of insurance beyond self-insurance.
@@
Guvenen et al. (2024) simulate a model with retirement, stochastic lifetimes, and altruism. The consumption response varies with the size and sign of shocks; the pass-through of the average negative permanent shock is between 0.29 and 0.46, depending on age, which is higher than ours but lower than BPP. They show that measuring pass-through as in BPP overestimates the true transmission rate by 1.5 times.
@@
We read these results as not too different from ours. Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow. Therefore, such shocks end up affecting consumption more than other shocks. The magnitude of our pass-through rate of permanent shocks is uniformly lower than its structural counterpart (consumption transmission in the data likely reflects multiple insurance sources, beyond self-insurance and taxes examined in these models), yet other empirical studies using the updated PSID, e.g., ABB and Arellano et al. (2023), find similar rates to ours.
Finally, we obtain back-of-the-envelope estimates of the welfare costs of tail income risk. Specifically, we simulate permanent income shocks from mixtures of normal distributions, we use those shocks to simulate consumption responses based on the quasi-reduced form consumption function in this paper, and we use power utility to calculate the discounted present value of a lifetime consumption stream associated with a given profile of income risk. Our baseline choice is tail risk similar to what we estimate in the data, while our counterfactual is Gaussian risk with similar variance. We then estimate the welfare costs of tail risk as the fraction of lifetime consumption a household facing Gaussian risk (counterfactual) would give up to avoid tail risk (baseline). We provide full details in appendix D.
Households facing large downside risk (ranging from a permanent shock 3 standard deviations below the mean at least once in their lifetime – 40% of households experience this – to a shock 5 standard deviations below the mean – a year’s long unemployment, 17% of households experience this at least once) would give up 4%-8% of lifetime consumption when income risk transmits linearly. This rises to 6%-12%, almost up to 1/8 of lifetime consumption, in the presence of nonlinear transmission. These estimates vary across the income and wealth distributions (appendix D), but the results remain in the same ballpark.
Although we calculate the welfare costs outside of a full structural model (among other things, taking wealth as given; hence ‘back-of-the-envelope’), our estimates seem to fall in the same ballpark as estimates from the structural papers described above. As a comparison, households in Busch and Ludwig (2024) sacrifice 0.5%-14% of lifetime consumption to live in a world with Gaussian risk, while households in Guvenen et al. (2024) sacrifice almost 22% of consumption to live in a world without any risk.